OpenStudy (anonymous):

@PaxPolaris What is the midpoint of a segment whose endpoints are (3, −1) and (−5, −3)?

OpenStudy (anonymous):

A. (4, −1) B. (−1, −2) C. (2, 1) D. (−2, 1)

OpenStudy (anonymous):

A. (4, −1) B. (−1, −2) C. (2, 1) D. (−2, 1)

OpenStudy (paxpolaris):

x coord: (3 + -5)/2 =-1 y coord (-1 + -3)/2 =-2 B. (-1,-2)

OpenStudy (anonymous):

What is the distance between the points (10, 2) and (14, 5)?

OpenStudy (anonymous):

A. 1 B. 3 C. 5 D. 10

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (anonymous):

Last one I promise. What is the midpoint of a segment whose endpoints are (3, 7) and (7, 3) A. (2, −2) B. (5, −5) C. (2, 2) D. (5, 5)

OpenStudy (paxpolaris):

 \[\sqrt{3^2+4^2}=5\]

OpenStudy (paxpolaris):

the average of 3 and 7 is 5 ... so...

OpenStudy (anonymous):

d?

OpenStudy (paxpolaris):

yes :)

OpenStudy (anonymous):

thank you so much!